OpenStudy (anonymous):

what is the value of x in the proportion (x minus1) over 5 equals (4x plus 2) over 35?

OpenStudy (anonymous):

\[\frac{ x-1 }{ 5 }=\frac{ 4x+2 }{ 35 }\]